### v3

# Exploring the Privacy-Energy Consumption Tradeoff for Split Federated Learning

Abstract
Split Federated Learning (SFL) has recently emerged as a promising distributed learning technology, leveraging the strengths of both federated and split learning. It emphasizes the advantages of rapid convergence while addressing privacy concerns. As a result, this innovation has received significant attention from both industry and academia. However, since the model is split at a specific layer, known as a cut layer, into both client-side and server-side models for the SFL, the choice of the cut layer in SFL can have a substantial impact on the energy consumption of clients and their privacy, as it influences the training burden and the output of the client-side models. In this article, we provide a comprehensive overview of the SFL process and thoroughly analyze energy consumption and privacy. This analysis considers the influence of various system parameters on the cut layer selection strategy. Additionally, we provide an illustrative example of the cut layer selection, aiming to minimize clients’ risk of reconstructing the raw data at the server while sustaining energy consumption within the required energy budget, which involves trade-offs. Finally, we address open challenges in this field. These directions represent promising avenues for future research and development.
I Introduction
Federated Learning (FL) fundamentally addresses the challenges associated with centralized learning by distributing the training process across multiple clients, enabling parallel processing. This approach also helps safeguard the privacy of raw data stored on clients by exchanging only model parameters. However, FL necessitates local training for each client, which can significantly burden clients with limited battery power and computational resources when dealing with large models like Deep Learning (DL). Split Learning (SL) has emerged as a solution to mitigate this problem. SL involves breaking down a full DL model into two sub-models, which can be trained both at a main server and across distributed clients. This approach alleviates the local training burden associated with FL while preserving data privacy. Nevertheless, SL introduces its own set of challenges, primarily related to training time overhead, owing to its relay-based training method. In this relay-based approach, only one client trains with the main server at any given time while other clients remain idle. This sequential training method leads to inefficient distributed processing and long training latency. To address this challenge, various strategies for parallelizing the training process of SL have been introduced [1]. Inspired by these efforts, Split Federated Learning, simply called SplitFed Learning (SFL), has been recently proposed as a novel approach that leverages the strengths of both FL and SL. Unlike SL, in SFL, all clients perform their local training in parallel while actively engaging with the main server and the federated server (fed server). In SFL, the fed server plays a pivotal role in aggregating the local model updates from clients using pre-defined aggregation techniques, such as FedAvg. This aggregation process occurs synchronously in each round of training. By introducing this additional aggregation server, SFL combines the advantages of both FL and SL seamlessly [2].
Figure 1: Workflow of SFL: In FL, the full model undergoes local training rather than just a subset. In SL, training processes do not occur in parallel. SFL seamlessly integrates the benefits of both FL and SL by addressing their respective limitations.
Despite the advantageous integration of SL and FL in SFL, the current SFL still presents several privacy concerns. Notably, the exchanged output of the client-side models, known as smashed data, and model updates between clients and servers have correlations with the raw data, leaving them susceptible to potential reconstruction attacks, where adversaries attempt to reconstruct the raw data from the correlated information. To tackle this growing concern, this article advocates for determining an appropriate split point between client-side and server-side models, specifically known as cut layer selection in the SFL, to mitigate the risk of reconstruction attacks while sustaining energy consumption within the required energy budget. This is particularly important for clients operating on limited battery power. It is worth noting that the selection of the cut layer significantly influences the outcome of the smashed data, impacting computational and communication overhead, as well as the level of privacy. For example, as discussed in [3], an ongoing study focused on the adaptive selection of cut layer, taking into account the varying computing and networking capabilities of clients. Nevertheless, prior research has not thoroughly analyzed both energy consumption and privacy in the context of cut layer selection, despite some studies focusing on empirical and analytical studies of reducing overall latency in SFL. This motivates us to provide an analysis of both energy consumption and privacy levels related to cut layer selection and a potential solution for developing novel cut layer selection techniques.
The contributions of this article can be summarized as follows:
  • We provide a comprehensive overview of the overall process of SFL along with an in-depth examination of how the choice of the cut layer impacts SFL in terms of both privacy and energy consumption.
  • Building upon this analysis, we propose a potential solution for optimizing cut layer selection, seeking to minimize the risk of reconstruction attacks while ensuring that energy consumption remains within the specified energy budget.
  • Finally, we shed light on several prospective avenues for future research and conclude the paper.
II Why is Cut Layer Selection Important?
II-A Background about SFL
The SFL framework consists of two servers such as i) a fed server and ii) a main server and multiple clients. The entire model is divided into two distinct sub-models: i) the client-side model and ii) the server-side model. The workflow of SFL is depicted in Fig. 1. Each client performs forward propagation on the client-side model using its dataset, passing the smashed data and corresponding label to the centralized main server (Step 1-2). Notably, the smashed data is the features extracted from its dataset via forward propagation using the client-side model. The main server performs forward propagation from the smashed data on the server-side model and back-propagation by calculating the loss between the true label and the predicted label, which can be done in parallel. Afterward, the server-side model is obtained by using FedAvg (Step 3-4) and the main server passes the gradient of the smashed data to each client for the client-side local model update (Step 5-6). Then, the fed server receives all clients’ updated client-side local models and aggregates them using FedAvg (Step 7-8). Finally, the fed server sends all clients the updated client-side global model, enabling synchronization of client-side models (Step 9). In summary, in the SFL approach, clients are responsible for training only a part of the model, which reduces their computational burden compared to traditional FL. Furthermore, SFL enhances the convergence speeds of SL by synchronizing client-side models through the fed server. Nonetheless, interacting with two servers in SFL introduces an additional communication overhead.
Figure 2: Example of reconstruction and membership inference attacks in SFL as a privacy concern: This highlights the tradeoff between the client model complexity and privacy. Note that A reconstruction attack is an attempt to reconstruct the original input data from the smashed data. Additionally, membership inference attacks are another type of attack, aiming to determine whether a given data record was part of the target model’s training dataset or not. (The random forest-based attack model is commonly employed in such scenarios.)
II-B A Brief Survey of SFL
As we discussed, SFL imposes significant communication overhead. This is primarily due to the transmission of smashed data, gradients and model updates between clients, the main server, and the fed server, indicating the need for improvements [4]. Nevertheless, there remains a scarcity of research focusing on the analytical modeling of SFL, in contrast to the abundance of empirical studies [5]. Notably, [6, 2] have contributed analytical models that investigate the impact of total model training time (i.e., latency) concerning the cut layer point. Correspondingly, based on such analysis, there have been various approaches to improve communication efficiency in SFL [4, 6, 1]. Specifically, [6] introduced an optimal cut layer selection based on latency analysis.
Moreover, In SFL, privacy concerns arise from the interactions between clients and servers, including the main and fed servers. During the forward phase, when clients transmit smashed data to the main server, this data becomes susceptible to reconstruction attacks , jeopardizing the privacy of the original information. Furthermore, the exchange of model updates with the fed server introduces another potential vulnerability to the raw data. To mitigate these risks, several privacy-preserving mechanisms can be employed. In a recent study [7] on SL, researchers conducted an empirical investigation into the impact of selecting the cut layer on reconstruction attacks during the forward phase. Their findings indicate that a greater depth of layers on the client side implies more non-linear functions, which compress the raw data by eliminating less informative features, thereby enhancing its resistance to reconstruction attacks. Similarly, the work of [8] also presented empirical studies on how the distance correlation between raw data and smashed data influences privacy leakage in SL about the selection of cut layer. Remarkably, most prior research has concentrated on analyzing and optimizing latency in SFL. Yet, there is a notable absence of studies addressing the impact of energy consumption concerning the cut layer selection in SFL. This aspect is crucial for estimating the associated costs for clients with limited battery power. Moreover, building on the insight of SL regarding privacy concerns, we will similarly conduct empirical studies of SFL, which can be extended to explore cut layer selection strategies that strike a balance between energy consumption and privacy level.
Figure 3: Proposed cut layer selection considering both energy consumption and privacy level.
II-C Cut Layer Selection Impact
II-C1 Energy Efficiency Perspective
In the context of SFL, energy consumption of clients occurs from both networking and computing perspectives. Specifically, from a computing perspective, clients engage in local training on their datasets using client-side models. From a networking perspective, there is energy expenditure associated with uplink transmission and downlink receiving between clients and the federated server as well as main server. Regarding downlink receiving energy consumption, it tends to be negligible due to the relatively small amount of energy required compared to uplink transmission power. In the case of uplink transmission energy consumption, energy is needed for transmitting smashed data to the main server. Additionally, each client must transmit its client side local model to the fed server for aggregation. It is worth noting that the cut layer selection strategy can have an impact on the size of the client-side model (i.e., model complexity). Consequently, as the size of the client-side model increases, energy consumption for both computation and model transmission to the fed server also increases. Therefore, to reduce energy consumption, a sophisticated management strategy for cut layer selection should be considered.
II-C2 Privacy Level Perspective
As shown in Fig. 2 , when we assume that the main server is honest but curious (HBC), it is still possible for the main server to reconstruct the original input data from the smashed data (specifically, the embedded features) received from clients. This process is referred to as a reconstruction attack. In this context, we can assess the privacy risks associated with this reconstruction attack by measuring the human perceptual similarity between the original and the reconstructed images, known as the structural similarity index (SSIM) [^1] This study concentrates on image data in the reconstruction attack, a prevalent focus in this area. Recently, the significance of text data privacy in this attack has also been recognized. In this case, metrics such as recovery rate, fluency, and attack accuracy are used to evaluate the privacy risk.. The SSIM formula compares two images by measuring three factors: luminance, contrast, and structure. It calculates the comparison as $\operatorname{SSIM}(x,y)=l(x,y)c(x,y)s(x,y)$, where $x$ and $y$ are the pixel values of the images being compared. The luminance comparison is denoted by $l$, contrast comparison by $c$, and structural comparison by $s$. This metric, which ranges from [0, 1], is widely used as a metric to measure the attack and defense performance [9], where the value $1$ denotes the most similar and vice versa. Fig. 2 (a)-(d) illustrates the relationship between client model complexity and privacy leakage. It highlights the effects of altering the cut layer on the quality of images reconstructed at the main server. Notably, as the cut layer index increases, indicating a rise in model complexity, the quality of the reconstructed image degrades. This degradation is interchangeable with enhanced privacy. The reason is that the complexity of the model deployed at the client’s end introduces advanced non-linearities in its output. This added intricacy can make it more challenging for adversaries to reverse-engineer and retrieve private input data. Moreover, Fig. 2 (a)-(d) also presents an illustrative example to practically demonstrate the application of noise addition, a fundamental strategy in achieving differential privacy. Here, our focus will center on one widely adopted noise addition mechanism: the Gaussian mechanism where it is instrumental in introducing carefully calibrated noise to sensitive data, effectively masking individual information while preserving the utility of the overall dataset. The Gaussian mechanism employs a Gaussian distribution for this purpose, providing distinct privacy-accuracy trade-offs. In this context, it is evident that introducing Gaussian noise to smashed data further degrades the quality of the reconstructed image, as indicated by the mean square error (MSE) representing the differences in pixels between the original and the reconstructed images. As the level of Gaussian noise rises, the MSE also increases, leading to a decrease in the SSIM in both shallow and deep cut layer scenarios. We also conduct another privacy assessment of the Membership Inference Attack (MIA), which is another type of attack on data privacy: determine whether a given data record was part of the target model’s training dataset or not. Here, the main server can still exploit these vulnerabilities to train a binary classifier and infer the membership information by taking the predicted value of the data sample obtained by querying the target model as input to the binary attack model. In this context, we can assess the privacy risks associated with this MIA by measuring the accuracy, which was widely used as a metric in MIA; the percentage of data records correctly inferred by the attack model in the whole given dataset. To train the attack model, we use transfer-inherit shadow learning. Fig. 2 (e) illustrates that MIA accuracy decreases with deeper cut layers and higher noise levels. Cutting at a shallow layer results in the main server retaining significant information about the entire target model, requiring less effort in the creation of a shadow client model. This makes the SFL vulnerable to MIA attacks, while the attack accuracy becomes more resilient to noise.
III Case Study: Cut Layer Selection
III-A An Illustrative Example
Based on the latency analysis of SFL in [2] and [6], and leveraging the energy consumption insights gained from FL studies [10], this section examines the overall energy consumption associated with SFL. Note that the analysis is based on the vanilla version of SFL rather than SFLv2/v3. However, the core concepts and intuitions are easily applicable to the higher versions. Similar to FL, as shown in Fig. 3, the total energy consumption $E$ of each client during SFL encompasses both i) computing and ii) networking energy. However, it requires customization to account for model splitting and interactions with both the fed server and the main server, distinctive features introduced in the SFL. Taking these considerations into account, to formulate the energy consumption model, consider a scenario with $K$ clients. Then, $K$ clients engage in the SFL framework for $g_{e}$ global epochs, representing the overall number of rounds necessary to achieve a specific training loss [11]. The client-side models synchronize at each global epoch, enabled by the fed server. In each global epoch, leveraging this synchronized client-side model, both clients and the main server conduct training for $l_{e}$ local epochs. Specifically, for simplicity, as in [2] and [6], considering homogeneous clients in a single local epoch, each client trains $D_{b}$ randomly sampled data items, commonly referred to as a mini-batch, from its dataset. By using transmission power $P_{t}$, they can communicate with the fed server and main server with total uplink (downlink) transmission rates $R_{1}^{U}$ and $R_{2}^{U}$ ($R_{1}^{D}$ and $R_{2}^{D}$), respectively. Consequently, the uplink (downlink) transmission rates for each client reduce to $\frac{R_{1}^{U}}{K}$ and $\frac{R_{2}^{U}}{K}$ ($\frac{R_{1}^{D}}{K}$ and $\frac{R_{2}^{D}}{K}$), respectively. Note that the entire model $W=[W_{C};W_{S}],$ where $W_{C}$, and $W_{S}$ are the client-side model, and the server-side model, respectively. As in [2] and [6], we also assume the same size of smashed data denoted as $q$ sent from clients to the main server. Let $|W|$ be the number of model parameters in the entire model $W$ where all model parameters have the same size $b$, and let $\alpha$ be the fraction of model parameters, which serves as the model cut layer point for SFL in $W_{C}$, where $|W_{C}|=\alpha|W|$ and $|W_{S}|=(1-\alpha)|W|$ [^2] As the depth of a DL model increases, there is a corresponding rise in both the number of learnable parameters and multiply-add floating-point operations [12]. Conversely, SFL divides the layers of a model and assigns only a fraction determined by $\alpha$ to a client, indicating that lower $\alpha$ values correspond to reduced client-side model complexity.. And, $T$ represents the time taken for one forward and backward propagation on the full model. In backward propagation, gradients of a constant size, denoted as $q^{\prime}$, are transmitted. Here, $P_{c}$ represents the power consumption necessary to train the full model $W$ at the client, which is calculated as the total energy consumption required to train the entire model, divided by the duration $T$. Therefore, in SFL, energy consumption during interaction with the main server, denoted as $E_{\text{main}}$, requires each client i) to proceed forward and backward propagation using $W_{C}$, incurring energy consumption $\alpha TP_{c}$ and ii) to transmit the smashed data (or receive the gradients) to (from) the main server per each sampled data item, incurring energy consumption $\frac{qK}{R_{2}^{U}}P_{t}$ (or $\frac{q^{\prime}K}{R_{2}^{D}}P_{r}$). Here, $P_{r}$ is the receiving power consumption at each client. In parallel, energy consumption for interaction with the fed server, denoted as $E_{\text{fed}}$, requires each client i) to send its local model to the fed server aggregation, incurring energy consumption $\frac{\alpha b|W|K}{R_{1}^{U}}P_{t}$, and ii) to receive the updated global model from the fed server, incurring energy consumption $\frac{\alpha b|W|K}{R_{1}^{D}}P_{r}$ during one global epoch. Finally, total energy consumption $E(\alpha)$ of each client for the SFL is given by
$E(\alpha)=g_{e}(l_{e}D_{b}E_{\text{main}}+E_{\text{fed}}).$ (1)
As previously discussed, under the assumption of an HBC main server, the privacy risks associated with the reconstruction attack can be evaluated by measuring the SSIM between the original input and the reconstructed images with respect to $\alpha$, denoted as $RS(\alpha)$. To evaluate the impact of cut layer selection in terms of $RS(\alpha)$ and $E(\alpha)$, we utilize a training-based adversarial inversion approach [13]. Note that the reconstruction model is implemented to reflect the inverted structure of the client-side model $W_{C}$ with transposed convolutional and Tanh activation layers. In our experiment, the overall classifier model $W$ for SFL is designed with four convolution blocks (Conv-BN-ReLU-Conv-BN-ReLU-Conv-BN-Max-ReLU) and one linear block. Linear block consists of two fully-connected layers and a Softmax function. The cut layer is selected between blocks to divide client and server-side models. For performance evaluation, the Fashion-MNIST dataset was employed [^3] https://github.com/zalandoresearch/fashion-mnist. The input image was resized to $32\times 32$ and the $D_{b}$ was 128. We used an Adam optimizer and the learning rate was set to 0.0002. The betas for Adam were set to (0.5, 0.999). We trained the classifier model for SFL up to 50 global training epochs and 75 local training epochs. The reconstruction model was trained up to 50 epochs under the same optimizer and settings as the classifier model. More detailed parameters are summarized at Table I.
TABLE I: Parameter Settings.
Parameters Values
$K$ 5
$W$ 31,484,464
$D_{b}$ 128
$q$, $q^{\prime}$ 491,520 bits, 491,520 bits
$b$ 32 bits
$T$ 0.00055 seconds
($R_{1}^{U},R_{2}^{U}$), ($R_{1}^{D}$, $R_{2}^{D}$) 200 Mbit/s, 100 Mbit/s
$P_{c}$, $P_{t}$, $P_{r}$ 4W, 0.2W, 0.2W
Figure 4: Reconstruction score and energy consumption with respect to the depth of cut layer: This represents the tradeoff between energy consumption and privacy.
As shown in Fig. 4, increasing the depth of the cut layer (i.e., increased complexity of the client-side model) leads to higher energy consumption ($E$), causing both $E_{\text{main}}$ and $E_{\text{fed}}$ to rise due to the increased computing and networking energy usage. Concurrently, RS consistently decreases with deeper cut layers. This phenomenon implies that the depth of the cut layer can be directly correlated with the complexity of the client-side model, potentially making attackers to reconstruct the image from the smashed data more difficult. Nevertheless, due to the lack of analytic understanding of RS, a data-driven approach can be applied to model RS through offline training using empirical measurements. In this context, we adopt regression-based modeling, which is one of the most widely used approaches, including mobile CPU property models (e.g., CPU power and temperature variation modeling). Consequently, we demonstrated that a convex model can effectively approximate the behavior of RS, yielding a Root Mean Square Error (RMSE) of only $0.0028$. As a result, RS can be estimated as $\operatorname{RS}(\alpha)$ = $0.76750.3597\alpha^{2}-0.7004\alpha+$ 0.7675. It should be noted that the estimated $RS(\alpha)$ may differ across applications; therefore, it should be pre-calculated in an offline manner prior to making any adjustments. Consequently, we can optimize the selection of the cut layer, denoted by $\alpha$, to minimize privacy leakage from reconstruction attacks, as measured by $RS(\alpha)$, while ensuring that energy consumption ($E$) remains within an acceptable energy budget ($E_{\text{req}}$). Thus, the problem can be formulated as follows:
[…]
Figure 5: Privacy gain of proposed scheme: It represents that optimal cut layer selection ($\alpha^{*}$=0.4201) achieves minimum privacy leakage (SSIM) while satisfying the required energy budget.
III-B Other Types of Control Variables
  • Computation/Power/Radio resource management: The computation resource, transmission power, and radio resource (i.e., bandwidth allocation over clients) of each client participating in SFL plays a pivotal role in optimizing performance. However, due to the early stage of this field, there has been a lack of rigorous analysis for SFL. Fundamentally, those factors are known to affect overall latency, denoted as $T_{i}$ in our analysis, which in turn impacts energy consumption within the field of FL [10].
  • Client and dataset management: The high mobility of clients can affect the wireless network conditions between clients and the main server and the fed server, respectively. Additionally, variations in the size of datasets across clients, characterized by their unbalanced, non-independent, and identically distributed (non-IID) nature, may lead to slower convergence speeds. Given these challenges, implementing strategies for dynamic client selection and effective dataset management—such as downsampling or augmenting datasets—within SFL can be beneficial in improving both accuracy and convergence speeds.
As we reviewed, the performance of SFL, concerning both energy efficiency and convergence speed, etc., is noteworthy. The heterogeneity in clients’ computing and networking capabilities can impact the speed of convergence due to the issue of stragglers, and this diversity also leads to differences in energy consumption among clients. Correspondingly, achieving an optimal balance concerning these multiple variables often requires substantial complexity in pursuing a global optimum. This necessitates adopting a modular design strategy or employing various heuristic methods to achieve sub-optimal solutions.
IV Open Challenges
IV-A Deep Reinforcement Learning
Considering the dynamic and heterogeneous nature of client networking and computing capabilities, deep reinforcement learning (DRL) shows promise in intelligently selecting the optimal cut layer. This aims to balance privacy, energy consumption, and convergence speed in the SFL, along with other essential control parameters, as discussed. Recent research by [15] illustrated how DRL is utilized to optimize cut layer selection for collaborative DL inference in multi-access computing (MEC) servers by efficiently offloading certain inference tasks from clients to the MEC server. The key challenges in applying DRL to address various SFL issues revolve around precisely defining the agent, environment, state, action, and reward. Additionally, selecting an appropriate DRL model and fine-tuning hyperparameters become pivotal for improving convergence speed, especially when dealing with complex and dynamic environments.
IV-B Privacy and Security Protection
As in [9], when clients and the main server hold their data and labels, respectively, attackers could be malicious data owners or eavesdroppers trying to intercept gradients shared from the main server to clients. The similarity in cut layer gradients among data samples can disclose their connection to labels, known as a label inference attack. Therefore, to mitigate risks, label inference attacks must be addressed by carefully choosing the cut layer and implementing differential privacy (DP). Moreover, determining the cut layer is also critical for maintaining security and integrity, particularly concerning model poisoning and data integrity. Specifically, the cut layer selection significantly influences the system’s susceptibility to model poisoning attacks. For instance, extensive client-side training elevates the risk of client-side poisoning, complicating the defense against distributed attacks. Thus, sophisticated aggregation techniques are required to counter model poisoning risks effectively. An inappropriately chosen cut layer can also hinder data authenticity and integrity validation, posing challenges to maintaining the model’s training integrity. The selection of the cut layer, therefore, demands careful consideration, balancing between minimizing security risks and maximizing the efficiency of SFL.
IV-C Lightweight Design
To enhance the efficiency of SFL, it is beneficial to consider the quantization approach. In the context of SFL, adopting a quantization approach can substantially reduce the size of both local model updates and intermediate data, which is useful particularly in resource-constrained clients. This reduction has several advantages, including reduced bandwidth usage, lower transmission energy, and reduced storage needs. Furthermore, quantization introduces inherent privacy guarantees. This is because reducing the precision of transmitted data inherently obfuscates the exact values, thereby adding an extra layer of protection for sensitive information, which is crucial in maintaining data privacy. Moreover, to mitigate the burden of training on resource-constrained clients with reduced energy consumption, the strategy of downsizing the resolution and the total volume of training data can be employed, considering different data distributions and acceptable accuracy levels. Additionally, selecting a lightweight DL model or employing model pruning techniques is adopted to facilitate a lightweight design for SFL.
V Conclusions
This article has provided an overview of the SFL, mainly focusing on its communication efficiency and privacy issues. By studying the impact of the cut layer selection on both energy consumption and privacy, we have provided a concrete example of an efficient cut layer selection to minimize the risk of reconstruction attacks within the required energy budget. Finally, we have suggested various other adjustable factors and highlighted some promising research directions within this field.
Biographies
Joohyung Lee (j17.lee@gachon.ac.kr) is currently an Associate Professor with the School of Computing, Gachon University and also a Visiting Fellow with the Department of Electrical and Computer Engineering, Princeton University. Before joining Gachon University, he was a Senior Engineer with Samsung Electronics From 2014 to 2017. His current research interests include federated learning, 6G networks, cloud/edge computing, smart grids, augmented reality, virtual reality, and network economics.
Mohamed Seif (mseif@princeton.edu) is currently a Post-Doctoral Research Associate with Princeton University, Princeton, NJ, USA. His current research interests include information theory, machine learning, and wireless communications.
Jungchan Cho (thinkai@gachon.ac.kr) is currently an Assistant Professor in the School of Computing at Gachon University, Seongnam, South Korea. From 2016 to 2019, he was a Senior Software Engineer at Samsung Electronics. His research interests include deep learning, computer vision, and machine learning.
H. Vincent Poor (poor@princeton.edu) is the Michael Henry Strater University Professor at Princeton University. His research interests are in the areas of information theory, machine learning and network science, and their applications in wireless networks, energy systems and related fields. Dr. Poor is a member of the U.S. National Academy of Engineering and the U.S. National Academy of Sciences. He received the IEEE Alexander Graham Bell Medal in 2017.